OpenStudy (anonymous):

write the equation 6^5=7776 in logarithmic form

OpenStudy (anonymous):

log base A= C base^C = A 6^5 = 7776 log (6) of 7776 = 5

OpenStudy (anonymous):

write the equation 243^1/5=3 in logarithmic form

OpenStudy (anonymous):

Same form! You try first :) I'll help if you try it

OpenStudy (anonymous):

write the equation (log1024) 16=2/5 in exponential form